OpenStudy (anonymous):

I need help with two math problems. I have tried working them out many times myself. and keep getting the wrong answer. System of equations: solve by addition. 1: .2x - .5y = -.4 -.3x + .4y = -.1 solution = (3,2) 2: 1/2x - 1/3y = 1 1/4x - 1/9y = 2/3 solution = (4,3) please explain what you are doing, I am only asking for help so i can pass my test.

OpenStudy (vlery):

all I can say is use guess and check sorry if it's not that helpful but that's what I do when idk what to do

OpenStudy (anonymous):

you said you tried solving it. can i see your solution so that we can find what you did wrong?

OpenStudy (anonymous):

these are my last attempts for these two problems

OpenStudy (dumbcow):

ok it seems the decimals threw you off a bit for #1 and you did more work than you need to why dont i just post a solution with explanation first of i would get rid of decimals by multiplying everything by 10 2x - 5y = -4 -3x +4y = -1 Now instead of adding them right away, you want to change the coefficients so that a variable is eliminated after adding Lets eliminate the "x" , 2 and 3 have common multiple of 6 multiply first equation by 3, 2nd equation by 2 6x - 15y = -12 -6x + 8y = -2 ----------------------- 0 -7y = -14 solve for y y = -14/-7 = 2

OpenStudy (dumbcow):

plugging y back in --> 2x - 5(2) = -4 2x -10 = -4 2x = 6 x = 3

OpenStudy (anonymous):

That makes sense to me.

OpenStudy (dumbcow):

good for #2 , you can also get rid of fractions by multiplying by a common denominator 4 and 9 have common multiple of 36 --> multiply everything by 36 18x - 12y = 36 9x - 4y = 24

OpenStudy (dumbcow):

try to do rest on your own

OpenStudy (dumbcow):

here is a video to help explain how to solve system of equations https://www.khanacademy.org/math/algebra2/systems_eq_ineq/systems_tutorial_precalc/v/systems-with-elimination-practice

OpenStudy (dumbcow):

gotta go, good luck :)

OpenStudy (anonymous):

thank you so much. i appreciate it. and I think i got it this time.